OpenStudy (anonymous):

medal given for help r2-8r-20 can you show me how to break it down to show the work it say factor the expression ..show all work-

OpenStudy (whpalmer4):

\[r^2-8r-20\] To factor this, multiply the coefficient of the first term by the coefficient of the last term: 1*-20 = -20 Now, find a pair of factors of -20 that add to the coefficient of the middle term, -8 What are two numbers that multiply together to give you -20, and add together to give you -8?

OpenStudy (anonymous):

ok (r2-10r+2r-20) (r(r-10+2r-20) (r2-10r+2r-20) (r+2)(r-10)) is this the correct breakdown?

OpenStudy (anonymous):

Answe (r+2)(r-10)

OpenStudy (whpalmer4):

Yes, that is, although you've got some extra and/or missing parentheses in a few spots. Factors obviously are -10, 2: \[r^2 - 10r + 2r -20\]\[(r^2-10r) + (2r-20)\]\[r(r-10) + 2(r-10)\]\[(r-10)(r+2)\] checking our work: \[(r-10)(r+2) = r(r+2)-10(r+2) = r^2 + 2r - 10r -20 = r^2-8r -20\checkmark\]

OpenStudy (anonymous):

TY

OpenStudy (anonymous):

Can you check this one -Did I break it down right? 8x4=36x3 4x3(2x)+4x3(-9) 4x3(2x-9)

OpenStudy (whpalmer4):

Assuming that your first line is supposed to be \(8x^4-36x^3\) then you did it correctly \(4x^3(2x-9)=2*4*x^3*x-9*4x^3 = 8x^4-36x^3\)

OpenStudy (anonymous):

have to put = in front of each breakdown ?

OpenStudy (whpalmer4):

No was just doing all the work on one line... \[4x^3(2x-9) = 2x*4x^3 - 9*4x^3\]\[2x*4x^3-9*4x^3 = 8x^4-36x^3\] same thing, but more wear and tear on my fingers and keyboard :-)

OpenStudy (anonymous):

ty

OpenStudy (whpalmer4):

you're welcome!